A worker wants to turn over a uniform 1.0x103N rectangular crate by pulling at 30°on one of its vertical sides. The floor is rough enough to prevent the crate from slipping. Calculate the pull that is needed to just start the crate to tip.


1
Expert's answer
2021-07-08T13:26:47-0400

Gives


"F=1.0\times10^3N"

"\theta=30°"

"W=F.dr=Fdrcos\theta"

"W=1.0\times10^3\times2\times cos30"


"W=1.0\times10^3\times2\times\frac{\sqrt3}{2}=1732J"